WendyMorales2007:

Um, I kinda 4got what this is XD

WendyMorales2007:

xxxFrancisco14xx:

76 TRUST I TOOK THE DIAGNOSTIC

xxxFrancisco14xx:

Can I get best response>

kimberlydereza:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @xxxFrancisco14xx Can I get best response> \(\color{#0cbb34}{\text{End of Quote}}\) U don't ask for it

xxxFrancisco14xx:

@kimberlydereza thx for telling me

Hoodmemes:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @xxxFrancisco14xx Can I get best response> \(\color{#0cbb34}{\text{End of Quote}}\) u lowkey wonder why u dont get medals its bc u ask for them/u also never explain how u got your answer the correct way.

WendyMorales2007:

thank you :)

xxxFrancisco14xx:

@WendyMorales2007 Ty for being nice

xxxFrancisco14xx:

lmk if you need more help

WendyMorales2007:

ok :)

xxxFrancisco14xx:

kk

jhonyy9:

9,13,43,55 55-13 = 42 43 -9 = 34 42 +34 = 76 76/2 = 38 @XioGonz your opinion pls about this ? ty

XioGonz:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 9,13,43,55 55-13 = 42 43 -9 = 34 42 +34 = 76 76/2 = 38 @XioGonz your opinion pls about this ? ty \(\color{#0cbb34}{\text{End of Quote}}\) You are Correct